The American Diabetes Association recommends a blood glucose reading of less than 130 for those with Type 2 diabetes. Blood glucose measures the amount of sugar in the blood. Below are the readings for February for a person recently diagnosed with Type 2 diabetes.

117 123 117 118 98 114 124 111 108 116
123 119 99 111 109 124 102 116 114 105
98 112 122 110 99 114 126 113

a. What is the arithmetic mean glucose reading? (Round your answer to 2 decimal places.)

Arithmetic mean

b. What is the median glucose reading? (Round your answer to 1 decimal place.)

Median

c. What is the modal glucose reading?

Modal
Answer #1

Ans. The data is given by

98 2 1g6 99 2 98 4 02 losT 108 lo 9 11 o 222 I1 112 113 13 5 los 1 108 L log Ilo I 11 2 112 13 114 3 116 2 Il? 2 118 119 1 122 1 123 2 2t 12 342 232 234 20 118 21 124 2 12 6 Total 28 112 246 248 26 31 5 2 2 3 2 5 27 20

(a)

Arithmetic mean = 3152/28 = 112.57

(b). Median = (14th term + 15thterm)/2

=(114+114)/2 = 114

(c). Modal = the value of x which have maximum frequency

= 114